A fox fleeing from a hunter encounters a 0.730 m tall fence and attempts to jump it. The fox jumps with an initial velocity of 7.10 m/s at an angle of 45.0°, beginning the jump 2.11 m from the fence. By how much does the fox clear the fence? Treat the fox as a particle.
1
Expert's answer
2012-02-07T14:16:43-0500
S_up = V^2*sin(45)^2/(2g)=1.28 m. (the highest point)
t= S_hor/(V*cos(45))=0.419 (time to reach the fence)

S_up=V*sin(45)*t-gt^2/2=1.243 m (point when the fox is above the hence)

The Fox will jump over the hence. t=0.419 sec